OpenStudy (anonymous):

On a science test, 2 points are deducted from a total of 100 points for each question answered incorrectly. Trevor scored an 84 on the test. The equation 100 – 2x = 84, where x represents the number of questions answered incorrectly, represents the situation. How many questions did Trevor answer incorrectly?

OpenStudy (chaise):

100-2x=84 -2x=84-100 -2x=-16 x=-16/-2 x=8 Do you need more help with rearranging equations?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

Subtract 100 from 84 then divide by 2. Then you will have your answer

OpenStudy (anonymous):

He got 16 wrong